A professor grades students on 5 quizzes, a project, and a final examination. Each quiz counts as 10% of the quiz grade. The project counts as 20% of the course grade. The final examination counts as 30% of the course grade. Tom has quiz scores of 80, 82, 88, 93 and 90. His project score is 96 and his final examination score is 95. Use the weighted mean formula to find Tom's average for the course *





1
Expert's answer
2021-12-06T02:19:57-0500

Consider the sum of all weights as


"5(10\% )+ 20\% + 30\% = 100\%"

Hence, the total course grade (TCG) is


"TCG =80(0.1)+82(0.1)+88(0.1)+93(0.1)"



"+90(0.1)+96(0.2)+95(0.3)=91"

Hence, the weighted mean (WM) is


"WM = \frac{91}{5(0.1)+0.2+0.2} = 91"